OpenStudy (anonymous):

What are the solutions to 3x^{2}=21x?

OpenStudy (anonymous):

Divide both sides by 3x (though don't forget 0 is also a solution)

OpenStudy (anonymous):

\[3x^{2} -21x =0.......................3x(x-7)=0................\therefore x=7,0 solutions\]

OpenStudy (anonymous):

\[3x^2=21x\]\[3x^2-21x=0\]\[x(3x-21)=0\]\[x=0\]or \[3x-21=0\]\[3x=21\]\[x=21/3\]\[x=7\]